# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 4.2 Learning Multi-Index Models with RCN
where we used the fact that $\mathbf{M}$ is psd. Which completes the proof of Claim 4.12. ∎
First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N{(\epsilon^{\prime}/k)^{\Omega(k)}}$.
Therefore, if $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, then by applying Claim 4.11 we have that with probability $1-\delta$ for any unit vector $\mathbf{v}\in W^{\perp}$
$\mathbf{v}^{\top}\widehat{\mathbf{U}}\mathbf{v}\leq\sum_{S\in\mathcal{S},\mathbf{u}\in U_{S}}(\mathbf{v}\cdot\mathbf{u})^{2}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\eta^{2}\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)\;.$
@@
Moreover, from Proposition 3.7 we have that if $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, then there exists a $\mathbf{v}\in\mathcal{E}$. Moreover, as $\mathbf{v}$ is a unit vector and $\left\|\mathbf{v}^{\perp W}\right\|\leq\epsilon$, we have that $\left\|\mathbf{v}^{W}\right\|\geq\sqrt{1-\epsilon^{2}}\geq 1-\epsilon$ which proves Proposition 4.10. ∎
Now we prove our main result for learning with RCN. Our proof similar to the proof of Theorem 3.3. The distinction is that we are going to use Proposition 4.10 which provides improved correlation when compared to Proposition 3.7.
@@
We show that Algorithm 5, with high probability, returns a hypothesis $h$ with 0-1 error at most $\tau+\mathrm{OPT}+\epsilon$. Let $f^{*}$ be an optimal classifier and denote by $W^{*}$ the subspace of its weight vectors. Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3d) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Denote by $T=2K$ the total number of iterations. Also let $\epsilon^{\prime}$ be the parameter set at Line 2 of Algorithm 5, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
@@
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ for all $t=1,\dots,T$. Using the fact that $N={(dm)^{Cm}2^{K^{C}}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{CK}\log(1/\delta)/(\sigma\alpha)^{C}}$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 5), we can apply Proposition 4.10 and conclude that, with probability $1-\delta$, there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t\in[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq 1-\epsilon\geq 1/2$. Thus, from Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-1/2$. After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T/2=K-T/2$. However, if $T$ is set to be $2K$ we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$ for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing by Claim 2.17, and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+2\epsilon$.
Sample and Computational Complexity: Note that the algorithm terminates in $O(K)$ iterations and at each iteration we draw $N={(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}}$ samples. Hence, the total sample size is ${(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}}$. Moreover we use at most $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time as all operations can be implemented in polynomial time in $N,d$ and $\lvert\mathcal{Y}\rvert$. ∎
@@ 5 Statistical Query Lower Bounds
Here we prove Theorems 1.5 and 1.7. First, in Section 5.1 we present SQ lower bounds about general classes of Multi-Index Models that do not satisfy Definition 1.2. Essentially showing that the algorithm of Section 3 is quantitatively optimal in the SQ model. Second, in Section 5.2, we present an SQ lower bound for the problem of learning multiclass linear classifiers under RCN.
@@ 5.1 SQ Lower Bounds for Learning Multi-Index Models
Here we show that if a Multi-Index Model class $\mathcal{F}$ does not satisfy Definition 3.2, any SQ algorithm that learns $\mathcal{F}$ to an accuracy comparable to that of Theorem 3.3 requires super-polynomial complexity. Specifically, we establish the following theorem:
Theorem 5.1.
(SQ Optimality of General Algorithm) Let $d,K,m\in\mathbb{Z}_{+}$ with $d^{c}\geq\max(m,K)$ for a sufficiently small absolute constant $c>0$, and let $\mathcal{Y}$ be a set of finite cardinality. Additionally, let $\tau,\mathrm{OPT}\in(0,1)$ with $\tau,\mathrm{OPT}\geq d^{-O(m)}$. Assume the existence of (i) a function $f:\mathbb{R}^{d}\mapsto\mathcal{Y}$ (ii) a random variable $y$ over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, (iii) a subspace $W$ of dimension at most $K$, and iv) a subspace $V$ such that:
@@
Let $\mathcal{F}$ denote the class of functions defined as =$\mathcal{F}=\{f(\mathbf{x}^{\perp U},\mathbf{U}\mathbf{x}^{U}):\mathrm{for\ all\ }\mathbf{U}\in\mathbb{R}^{k\times d}:\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{k}\}$, where $k=\dim(U)$. Then, any SQ algorithm that learns $\mathcal{F}$ up to an error of smaller than $\tau-2\mathrm{OPT}$ with probability at least $2/3$ requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(m)}/|\mathcal{Y}|\right)$.
@@
Let $\mathcal{Q}\subseteq\mathbb{R}^{d\times K^{\prime}}$ be the set of matrices from C.4 for which it holds that $\lvert\mathcal{Q}\rvert=2^{\Omega(d^{c})},\mathbf{Q}\mathbf{Q}^{\top}=\mathbf{I}_{K^{\prime}}$ and $\left\|\mathbf{Q}\mathbf{P}^{\top}\right\|_{F}\leq O(d^{2c-1+2\alpha})$ for all $\mathbf{Q},\mathbf{P}\in\mathcal{Q}$, where $\alpha=\log K^{\prime}/\log d$. For each $\mathbf{Q}\in\mathcal{Q}$, we define the function $f_{\mathbf{Q}}(\mathbf{x}):=f(\mathbf{x}^{\perp U},\mathbf{Q}\mathbf{x}^{U})$ along with the coupled random variables $y_{\mathbf{Q}}$ satisfying $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i|\mathbf{x}=\mathbf{z}]=\operatorname*{\mathbf{Pr}}[y=i|\mathbf{x}=\mathbf{z}^{\perp U}+\mathbf{Q}\,\mathbf{z}^{U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$. We consider the following class of functions $\mathcal{F}_{\mathcal{Q}}=\{f_{\mathbf{Q}}(\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}$ and note that by definition $\mathcal{F}_{\mathcal{Q}}\subseteq\mathcal{F}$.
For each $\mathbf{Q}\in\mathcal{Q}$, let $D_{\mathbf{Q}}$ be the joint distribution of $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Denote by $\mathcal{D}=\{D_{\mathbf{Q}}:\mathbf{Q}\in\mathcal{Q}\}$ the class of these distributions. Denote by $D_{0}$ the distribution of pairs $(\mathbf{x},y_{0})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{0}$ a random variable supported on $\mathcal{Y}$ that depends only on $W^{V}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[y=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$. Denote by $y_{i}(\mathbf{x}^{\perp U})=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}]$.
We first show that for all $D_{\mathbf{P}},D_{\mathbf{Q}}\in\mathcal{D}$ with $\mathbf{P}\neq\mathbf{Q}$, the correlation $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ is small. Using the law of total expectation, we have that
@@
Considering each term separately, by C.5, we have that
@@
$\leq\sum_{t=m}^{\infty}\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{t}\operatorname*{\mathbf{E}}[\mathds{1}(y=i)^{2}\mid\mathbf{x}^{\perp U}]\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]\;,$
where we used Parseval’s identity and Assumption (3) of the theorem statement. Therefore, we have that
$\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ $\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\frac{\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]}{y_{i}(\mathbf{x}^{\perp U})}\right]\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}|\mathcal{Y}|\;.$
Hence, $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})=d^{-\Omega(m)}\lvert\mathcal{Y}\rvert$. Similarly, we can show that $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{P}})\leq\lvert\mathcal{Y}\rvert$ for all $D_{\mathbf{P}}\in\mathcal{D}$. Therefore, we have that $\mathcal{D}$ is $(d^{-\Omega(m)}\lvert\mathcal{Y}\rvert,\lvert\mathcal{Y}\rvert)$-correlated relative to $D_{0}$ (see Definition C.1), and thus from C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D_{0})$ with probability at least $2/3$ requires $2^{d^{\Omega(1)}}$ queries to the $\text{VSTAT}(d^{\Omega(m)}/\lvert\mathcal{Y}\rvert)$ oracle.

### v2

@@ 4.2 Learning Multi-Index Models with RCN
where we used the fact that $\mathbf{M}$ is a PSD matrix. This completes the proof of Claim 4.12. ∎
First note that by the definition of an approximating partition (Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ and $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$. Therefore, by the union bound and Hoeffding’s inequality, it holds that if $N\geq(k/\epsilon^{\prime})^{Ck}\log(\lvert\mathcal{S}\rvert/\delta)=(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N(\epsilon^{\prime}/k)^{\Omega(k)}$.
Therefore, if $N\geq(dm)^{Cm}(k/\epsilon^{\prime})^{Ck}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}$ for a sufficiently large universal constant $C>0$, then by applying Claim 4.11 we have that the following: with probability $1-\delta$, for any unit vector $\mathbf{v}\in W^{\perp}$
$\mathbf{v}^{\top}\widehat{\mathbf{U}}\mathbf{v}\leq\sum_{S\in\mathcal{S},\mathbf{u}\in U_{S}}(\mathbf{v}\cdot\mathbf{u})^{2}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\eta^{2}\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)\;,$
@@
Moreover, by Proposition 3.7 we have that if $N\geq(dm)^{Cm}(k/\epsilon^{\prime})^{Ck}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}$ for a sufficiently large universal constant $C>0$, then there exists a $\mathbf{v}\in\mathcal{E}$. Moreover, as $\mathbf{v}$ is a unit vector and $\left\|\mathbf{v}^{\perp W}\right\|\leq\epsilon$, we have that $\left\|\mathbf{v}^{W}\right\|\geq\sqrt{1-\epsilon^{2}}\geq 1-\epsilon$ which proves Proposition 4.10. ∎
Now we are ready to prove Theorem 4.9. Our proof is similar to the proof of Theorem 3.3. The difference is that we are going to use Proposition 4.10 which provides improved correlation when compared to Proposition 3.7.
@@
We show that Algorithm 5, with high probability, returns a hypothesis $h$ with 0-1 error at most $\tau+\mathrm{OPT}+\epsilon$. Let $f^{*}$ be an optimal classifier and denote by $W^{*}$ the subspace of its weight vectors. Let $L_{t}$ be the set of vectors maintained by the algorithm (Line 3d) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Denote by $T=2K$ the total number of iterations. Also let $\epsilon^{\prime}$ be the parameter set at Line 2 of Algorithm 5, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
@@
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ for all $t=1,\dots,T$. Using the fact that $N=(dm)^{Cm}2^{K^{C}}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{CK}\log(1/\delta)/(\sigma\alpha)^{C}$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 5), we can apply Proposition 4.10 and conclude that, with probability $1-\delta$, there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and unit vectors $\mathbf{w}^{(t)}\in W^{*}$ for $t\in[T]$ such that $\mathbf{w}^{(t)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq 1-\epsilon\geq 1/2$. Thus, from Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-1/2$. After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T/2=K-T/2$. However, if $T$ is set to be $2K$ we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$ for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing (see Claim 2.17), and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+2\epsilon$.
Sample and Computational Complexity: Note that the algorithm terminates in $O(K)$ iterations and at each iteration we draw $N=(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}$ samples. Hence, the total sample size is $(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}$. Moreover we use at most $\mathrm{poly}(N)$ time as all operations can be implemented in polynomial time in $N,d$ and $\lvert\mathcal{Y}\rvert$. ∎
@@ 5 Statistical Query Lower Bounds
In this section we prove Theorems 1.6 and 1.9. First, in Section 5.1 we present SQ lower bounds for general classes of Multi-Index Models that do not satisfy Definition 1.2. Essentially showing that the algorithm of Section 3 is quantitatively optimal in the SQ model. Second, in Section 5.2, we present an SQ lower bound for the problem of learning multiclass linear classifiers under RCN.
Before presenting our hardness results, we first provide a short introduction to the SQ model.
Basics on SQ Model
SQ algorithms are a broad class of algorithms that, instead of having direct access to samples, are allowed to query expectations of bounded functions of the distribution.
Definition 5.1 (SQ algorithms).
Let $D$ be a distribution on $\mathcal{X}$. A statistical query is a bounded function $q:\mathcal{X}\to[-1,1]$. For $u>0$, the $\mathrm{VSTAT}(u)$ oracle responds to the query $q$ with a value $v$ such that $|v-\operatorname*{\mathbf{E}}_{\mathbf{x}\sim D}[q(\mathbf{x})]|\leq\tau$, where $\tau=\max(1/u,\sqrt{\mathrm{Var}_{\mathbf{x}\sim D}[q(\mathbf{x})]/u})$. We call $\tau$ the tolerance of the statistical query. A Statistical Query algorithm is an algorithm whose objective is to learn some information about an unknown distribution $D$ by making adaptive calls to the corresponding oracle.
The SQ model was introduced in [Kea98] as a natural restriction of the PAC model [Val84]. Subsequently, the model has been extensively studied in a range of contexts, see, e.g., [Fel16]. The class of SQ algorithms is broad and captures a range of known supervised learning algorithms. More broadly, several known algorithmic techniques in machine learning are known to be implementable using SQs (see, e.g., [FGR+13, FGV17]).
Throughout this section, we leverage a number of previously known facts about the SQ model that allow us to establish hardness. We refer the reader to Appendix C for their formal statements.
@@ 5.1 SQ Lower Bounds for Learning Multi-Index Models
We show that if a Multi-Index Model class $\mathcal{F}$ does not satisfy Definition 3.2, any SQ algorithm that learns $\mathcal{F}$ to an accuracy comparable to that of Theorem 3.3 requires super-polynomial complexity. Specifically, we establish the following theorem:
Theorem 5.2 (SQ Optimality of General Algorithm).
Let $d,K,m\in\mathbb{Z}_{+}$ with $d^{c}\geq\max(m,K)$ for a sufficiently small absolute constant $c>0$, and let $\mathcal{Y}$ be a set of finite cardinality. Additionally, let $\tau,\mathrm{OPT}\in(0,1)$ with $\tau,\mathrm{OPT}\geq d^{-O(m)}$. Assume the existence of (i) a function $f:\mathbb{R}^{d}\mapsto\mathcal{Y}$ (ii) a random variable $(\mathbf{x},y)$ over $\mathbb{R}^{d}\times\mathcal{Y}$ with $\mathbf{x}$-marginal the standard Gaussian such that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ (iii) a subspace $W$ of dimension at most $K$, and iv) a subspace $V$ such that:
@@
Let $\mathcal{F}$ denote the class of MIMs defined as =$\mathcal{F}=\{f(\mathbf{x}^{\perp U},\mathbf{U}\mathbf{x}^{U}):\mathrm{for\ all\ }\mathbf{U}\in\mathbb{R}^{k\times d}:\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{k}\}$, where $k=\dim(U)$. Then, any SQ algorithm that learns $\mathcal{F}$ in the presence of $\mathrm{OPT}$ adversarial label noise up to an error of smaller than $\tau-2\mathrm{OPT}$ with probability at least $2/3$ requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(m)}/|\mathcal{Y}|\right)$.
@@
Let $\mathcal{Q}\subseteq\mathbb{R}^{d\times K^{\prime}}$ be the set of matrices from C.4 for which it holds that $\lvert\mathcal{Q}\rvert=2^{\Omega(d^{c})},\mathbf{Q}\mathbf{Q}^{\top}=\mathbf{I}_{K^{\prime}}$ and $\left\|\mathbf{Q}\mathbf{P}^{\top}\right\|_{F}\leq O(d^{2c-1+2\alpha})$ for all $\mathbf{Q},\mathbf{P}\in\mathcal{Q}$, where $\alpha=\log K^{\prime}/\log d$. For each $\mathbf{Q}\in\mathcal{Q}$, we define the function $f_{\mathbf{Q}}(\mathbf{x}):=f(\mathbf{x}^{\perp U},\mathbf{Q}\mathbf{x}^{U})$ along with the coupled random variables $y_{\mathbf{Q}}$ satisfying $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i|\mathbf{x}=\mathbf{z}]=\operatorname*{\mathbf{Pr}}[y=i|\mathbf{x}=\mathbf{z}^{\perp U}+\mathbf{Q}\,\mathbf{z}^{U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$. We consider the class of functions $\mathcal{F}_{\mathcal{Q}}=\{f_{\mathbf{Q}}(\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}$ and note that by definition $\mathcal{F}_{\mathcal{Q}}\subseteq\mathcal{F}$.
For each $\mathbf{Q}\in\mathcal{Q}$, let $D_{\mathbf{Q}}$ be the joint distribution of $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Denote by $\mathcal{D}=\{D_{\mathbf{Q}}:\mathbf{Q}\in\mathcal{Q}\}$ the class of these distributions. Denote by $D_{0}$ the distribution of pairs $(\mathbf{x},y_{0})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{0}$ a random variable supported on $\mathcal{Y}$ that depends only on $W^{V}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y_{0})\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[y=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$. Denote by $y_{i}(\mathbf{x}^{\perp U})=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}]$.
We show that is SQ-Hard to determinate whether we have access to distribution $D_{0}$ or to a distribution $D\in\mathcal{D}$. To this end, we show that for all $D_{\mathbf{P}},D_{\mathbf{Q}}\in\mathcal{D}$ with $\mathbf{P}\neq\mathbf{Q}$, the correlation $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ (see Definition C.1) is small. Using the law of total expectation, we have that
@@
Considering each term separately, by C.5 and Assumption (3) of the theorem statement, we have that
@@
$\leq\sum_{t=m}^{\infty}\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{t}(y^{[t]})^{2}\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\sum_{t=m}^{\infty}(y^{[t]})^{2}\;,$
where $y^{[t]}$ is the degree-$t$ Hermite part of the random variable $\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}$ and that $\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}\leq 1$. Furthermore, using Parseval’s identity we get that $\sum_{t=m}^{\infty}(y^{[t]})^{2}\leq\operatorname*{\mathbf{E}}[\mathds{1}(y=i)^{2}\mid\mathbf{x}^{\perp U}]=\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]$. Therefore, we have that
$\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ $\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\frac{\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]}{y_{i}(\mathbf{x}^{\perp U})}\right]=\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}|\mathcal{Y}|\;,$
where we used that $\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[({\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]})/{y_{i}(\mathbf{x}^{\perp U})}\right]=1$, by the definition of $y_{i}(\mathbf{x}^{\perp U})$. Hence, $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})=d^{-\Omega(m)}\lvert\mathcal{Y}\rvert$. Similarly, we can show that $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{P}})\leq\lvert\mathcal{Y}\rvert$ for all $D_{\mathbf{P}}\in\mathcal{D}$. Therefore, we have that $\mathcal{D}$ is $(d^{-\Omega(m)}\lvert\mathcal{Y}\rvert,\lvert\mathcal{Y}\rvert)$-correlated relative to $D_{0}$ (see Definition C.1), and thus from C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D_{0})$ with probability at least $2/3$ requires $2^{d^{\Omega(1)}}$ queries to the $\text{VSTAT}(d^{\Omega(m)}/\lvert\mathcal{Y}\rvert)$ oracle.